OpenStudy (anonymous):

Find the solution set 2w+1-(5-w)=3(w-7)

OpenStudy (anonymous):

Distribute that - on the left side and distribute that 3 on the right side \[2w+1-5+w=3w-21\] Then combine terms on the right\[3w-4=3w-21\] Cancel the 3w on each side and you get the impossible \[-4=-21\] When you get an impossible statement like this the solution set is "No solution" or "the empty set"

OpenStudy (anonymous):

ahhhhhh okay :)

OpenStudy (anonymous):

one more question

OpenStudy (anonymous):

x^2-2x=-10 why is this equation, doesnt have solution????